OpenStudy (anonymous):

The area of a rectangle is 65 ft2 , and the length of the rectangle is 3 ft less than twice the width. Find the dimensions of the rectangle.

OpenStudy (kash_thesmartguy):

l=(w*2)-3 65=l*w Try using those formulas to calculate W and then find the measures

OpenStudy (kash_thesmartguy):

@dharmon14